OpenStudy (anonymous):

Why is tangent's period 180 degrees?

OpenStudy (anonymous):

I can't say why but how to find i can say that There's no splitting required here. The tangent is the sine, divided by the cosine. Looking at the unit circle, we see that the point for 180 degrees is all the way on the left, at (-1, 0). That gives us a cosine of -1 and a sine of 0. 0 divided by -1 = 0 So the tangent of 180 degrees is 0.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

It's because sin(x+180) = - sinx and cos(x+180) = -cosx, so tan(x+180) = sin(x+180)/cos(x+180) = -sinx/-cosx=sinx/cosx=tanx

OpenStudy (amistre64):

becasue every 180 degrees it has the same value

OpenStudy (amistre64):



OpenStudy (amistre64):

bad mouse!! bad!!